OpenStudy (anonymous):

given that cosA=3/5 and sinB=-15/17, with A in Q1 and B in Q4, use an appropriate sum formula to compute the exact value of cos(a+b)

ganeshie8 (ganeshie8):

\(\cos (A + B) = \cos (A) \cos (B) - \sin (A) \sin (B)\)

ganeshie8 (ganeshie8):

find watever u dont knw and plug them in above formula

OpenStudy (anonymous):

i realize that's the formula, but how do i find cos b and sin a?

ganeshie8 (ganeshie8):

use below identity : \(\sin^2\theta + \cos^2\theta = 1\)

ganeshie8 (ganeshie8):

you know cosA=3/5, can u find sinA ?

ganeshie8 (ganeshie8):

\(\large \sin^2A + \cos^2A = 1\)

OpenStudy (anonymous):

gotcha

ganeshie8 (ganeshie8):

\(\large \sin^2A + (\frac{3}{5})^2 = 1\)

OpenStudy (anonymous):

Think it just clicked.

ganeshie8 (ganeshie8):

solve \(\sin A\)

ganeshie8 (ganeshie8):

good :) but there is a trick, u need to fix sign based on the quadrant

OpenStudy (anonymous):

so since it's cos if its quadrant 1 or 4 its positive, 2 or 3 its negative?

ganeshie8 (ganeshie8):

\(\large \sin^2A + (\frac{3}{5})^2 = 1 \) \(\large \sin^2A = 1 - \frac{9}{25} \) \(\large \sin^2A = \frac{16}{25} \) \(\large \sin A = \pm \frac{4}{5} \)

ganeshie8 (ganeshie8):

since \(A\) is first Quadrant, "sin" is positive, so : \(\large \sin A = + \frac{4}{5} \)

OpenStudy (anonymous):

got it, thanks so much!

ganeshie8 (ganeshie8):

for cosB , you should take positive value ok cuz B is in fourth Quadrant

ganeshie8 (ganeshie8):

:) u wlc !

OpenStudy (anonymous):

got cos(a+b)=84/85. Strangest answer...